OpenStudy (anonymous):

A rectangle has width 22 in and length 18 in. Find the length of another rectangle of equal area, if the width is 20 in.

OpenStudy (anonymous):

area1 = 22*18 = 396 area2 = 20*x = 396. So, x = 396/20 = 19.8.